# Untangling Lariats: Subgradient Following of Variationally Penalized Objectives

### v1

@@ 6 High-Order Variational Penalties
To illustrate the power of the convolution-based representaion, we provide a specific derivation for variational penalties which correspond to second order finite differences. To do so we set $\boldsymbol{\tilde{\tau}}=(2,-1)$. Since $f_{i}(\cdot)$ depends only on $x_{i-1}$ and $x_{i}$, we can represent the subgradient sets as,
$S_{i}=\{\mathsf{a}_{i}(\mathsf{x}_{i-1},\mathsf{x}_{i})\;;\;\mathsf{a}_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i})\}~{}\mbox{ and }~{}S_{i+\frac{1}{2}}=\{\mathsf{a}_{i}^{\prime}(\mathsf{x}_{i},\mathsf{x}_{i+1})\;;\;\mathsf{a}_{i+1}^{\prime}(\mathsf{x}_{i},\mathsf{x}_{i+1})\}~{}.$
Assume this holds true for $S_{i+1}~{}\mbox{ i.e. }~{}S_{i+1}=\{\mathsf{a}_{i+1}(\mathsf{x}_{i},\mathsf{x}_{i+1})\;;\;\mathsf{a}_{i+2}(\mathsf{x}_{i},\mathsf{x}_{i+1})\}~{}.$ The definition of the recursive map for $i+\frac{1}{2}$ implies that,
$\mathsf{a}_{i}^{\prime}(\mathsf{x}_{i},\mathsf{x}_{i+1})=2\mathsf{a}_{i+1}(\mathsf{x}_{i},\mathsf{x}_{i+1})-\mathsf{a}_{i+2}(\mathsf{x}_{i},\mathsf{x}_{i+1})+\mathsf{x}_{i+1}-y_{i+1}~{}\mbox{ and }~{}\mathsf{a}_{i+1}^{\prime}(\cdot)=\mathsf{a}_{i+1}(\cdot)~{}.$
To obtain the recursive map $S_{i+\frac{1}{2}}\to S_{i}$, we need to solve for $\mathsf{x}_{i+1}$ from the following equation
$2\mathsf{x}_{i}-\mathsf{x}_{i-1}-\mathsf{x}_{i+1}\in\partial g_{i}^{*}(\mathsf{a}_{i}^{\prime}(\mathsf{x}_{i},\mathsf{x}_{i+1}))~{}.$
@@
$T_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i})=\big{[}2\mathsf{x}_{i}-\mathsf{x}_{i-1}\big{]}_{z_{i+1}^{-}(\mathsf{x}_{i})}^{z_{i+1}^{+}(\mathsf{x}_{i})}~{},$
where $z_{i+1}^{\pm}(\mathsf{x}_{i})$ denotes the boundary values such that $\mathsf{a}_{i}^{\prime}(\mathsf{x}_{i},z_{i+1}^{\pm}(\mathsf{x}_{i}))=\pm 1~{}.$ Using this form of $T$ back in $\mathsf{x}_{i+1}$, we obtain
@@
$\mathsf{a}_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i})=\mathsf{a}_{i+1}^{\prime}(\mathsf{x}_{i},T_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i}))~{}.$
@@
Here, $\mathsf{a}_{i}(\cdot)$ is a piecewise linear manifold in $\mathbb{R}^{2}$. We store triplets of the form $\{v_{j}^{1},v_{j}^{2},\alpha_{j}\}$ for vertices of each facets defining $\mathsf{a}_{i}$. We conjecture that the number of facets grows as $\mathcal{O}(n^{2})$ but leave formal analysis of runtime as future research. We empirically observed that the description of $\mathsf{a}_{i}$ is typically linear in the length of the input sequence. Alas, a disadvantage of the exact subgradient following described here is the need to maintain a data structure for high dimensional objects. We describe an alternative approximation yielding a simple iterative procedure for the problem in Appendix D.
@@
We next demonstrate the effectiveness of exact subgradient following over the approximation-based method. To do so, we generated synthetic data of varying sequence length similar to previous section. We sampled $(x_{i+2}-x_{i+1})-(x_{i+1}-x_{i})$ from $\mathcal{D}$ and then constructed $\mathbf{y}$ from optimality conditions. Throughout the experiment, the sparsity level was set to $50\%$. In each experiment, we ran the approximation-based algorithm for $100$ iterations. The results are summarized in Figure 7. The advantage of exact subgradient following over approximation both in terms of runtime and optimality is apparent. The exact subgradient following depends rather weakly on the problem size whereas the quality of the solution as well run time quickly deteriorates even for short sequences.
7 Sub-Discussion
@@
$\lambda_{i}^{\mathsf{R}}\big{[}x_{i+1}-x_{i}-\epsilon_{i}^{\mathsf{R}}\big{]}_{+}\,+\,\lambda_{i}^{\mathsf{L}}\big{[}x_{i}-x_{i+1}-\epsilon_{i}^{\mathsf{L}}\big{]}_{+}~{}.$
Bu choosing $\lambda_{i}^{\mathsf{R}}=\lambda_{i}^{\mathsf{L}}=\lambda_{i}$ and $\epsilon_{i}^{\mathsf{R}}=\epsilon_{i}^{\mathsf{L}}=0$ we obtain the fused lasso’s penalty. To encompass this version we simply need to compute $\partial g_{i}^{*}$ from which we derive a new mapping $\mathsf{x}\mapsto\mathsf{a}_{i}(\mathsf{x})$. The subgradient procedure uses a similar form for the mapping given in Eq. 7 though the introduction of insensitivity regions to the left and/or the right of the loci where $x_{i}=x_{i+1}$ would introduce two addition inflection points. The resulting procedure for constructing the subgradient would require at most $\mathcal{O}(n\log n)$ time.
We limited our discussion to sequential penalties which can described as a chain of local connectivities $x_{1}\leftrightarrow x_{2}\leftrightarrow\ldots x_{i}\leftrightarrow x_{i+1}\ldots$ forming a left-to-right dependency graph. Kolmogorov et. al [12] examined a setting where the total variation penalties form a tree. Let us denote the variation dependency graph by $G=(V,E)$ where $V=[n]$ and $(i,j)\in E$ if there exists a penalty term for the variables $x_{i}$ and $x_{j}$. That is, the penalized problem takes the form of,
@@
$U$ $=$ $\max_{\|\boldsymbol{\alpha}\|_{\infty}\leq 1}\,\min_{{\mathbf{x}}}\>\frac{1}{2}\sum_{i=1}^{n}(x_{i}-y_{i})^{2}+\sum_{(i,j)\in E}\alpha_{i,j}(x_{i}-x_{j})~{}.$
@@
[Uncaptioned image][Uncaptioned image]
This work opens numerous directions for applied research. We would like here to sketch a single domain of potential applications of trajectory smoothing. Consider the task of mimicking cursively written notes such as the word foundation depicted on the top right. While the writing is clearly intelligible the letters are not of the same size and the word’s baseline drifts up and down. This characteristics is common to all cursively written texts by humans, and by some macaque monkeys. However, a synthetically generated text by a brain-controlled prosthetic would benefit from a low-jerk and smooth motor control of the electromechanical apparatus in charge of moving the artificial limb. To this end, a dynamical system such as the spring model of Hollerbach [9] can be used to approximate the noisy observations while inflicting penalty on high-jerk motions using the tools introduced in Sec. 6. The result would resemble the text depicted on the bottom right which was generated by a context-dependent dynamic programming.
@@ Acknowledgments
We would like to thanks Tomer Koren and Yishay Mansour for fruitful discussions. The short discussion above is motivated by conversations with the late Prof. Tali Tishby who is missed dearly. This manuscript was written without the usage of AIML tools.
Appendix A Improving Runtime for Fused Barriers using Skiplist
@@
We store each $\mathsf{a}_{i}(\cdot)$ as a linked-list data structure of implicit change points (vertices) and explicit linear intervals connecting two consecutive vertices. A vertex $(v_{j},\mathsf{a}_{i}(v_{j}))=(v_{j},\alpha_{j})$ represents a point where $\mathsf{a}_{i}(\cdot)$ changes slope. Edge $(d_{j},s_{j})$ stores the length and slope of the line segment between $v_{j}$ and $v_{j+1}$,
$d_{j}=v_{j+1}-v_{j}\,;\,s_{j}=(\alpha_{j+1}-\alpha_{j})/d_{j}~{}.$
@@
This representation allows lazy evaluation in a similar fashion to the fused lasso. The Kvetsh step does not change the lengths of intervals except for the new segment encapsulating the zero crossing. Since the length of all other line segments remain intact this edge-based representation facilitates a stepping stone for an efficient implementation.
@@
By construction of $\mathsf{a}_{i+1}$ we know that $\exists j$ for which $v_{j}=z_{i+1}-\lambda_{i+1}$ and $\mathsf{a}_{i+1}(z_{i+1}-\lambda_{i+1})=0$. This implies that in an explicit representation of $\mathsf{a}^{\prime}_{i}(\cdot)$ we would need to update vertex $j$, $(z_{i+1}-\lambda_{i+1},0)\mapsto(z_{i+1}-\lambda_{i+1},z_{i+1}-\lambda_{i+1}-y_{i+1})$. In the explicit yet lazy representation of the ensuing edge, the slope $s_{j}$ of the segment $[v_{j},v_{j+1}]$ becomes $s_{j}+n-i$. Next we find the zero crossing point of $\mathsf{a}^{\prime}_{i}(\cdot)$ by traversing the list of edges in the direction towards zero and locating the index $k$ for which $\alpha_{k}\leq 0$ and $\alpha_{k+1}>0$. In case such an index does not exist the zero of $\mathsf{a}^{\prime}_{i}(\cdot)$ is right of its right-most vertex. Once we identify the encapsulating interval, we calculate $z_{i}$ by solving a linear equation from $(v_{k},\alpha_{k})$ and $(v_{k+1},\alpha_{k+1})$. We next Kvetsh the list by inserting two vertices between $k$ and $k+1$ whose content is $(z_{i}-\lambda_{i},0)$ and $(z_{i}+\lambda_{i},0)$ and updating $d_{k}$ accordingly. This sequential traversal of the list in the worst-case would require linear time and thus total run time would remain intact at $\mathcal{O}(n^{2})$.
@@
In order to facilitate a more efficient procedure for finding zero crossing. We make use of a skip-list data structure [14]. A skip-list is an extension of linked list with multiple layers and with which the expected search time is $\mathcal{O}(\log{n})$. The skip-list has ideally in our setting $\mathcal{O}(\log{n})$ layers where layer $k$ contains vertices whose index modulo $2^{k}$ is zero. The search procedure starts from highest layer and goes down once we reach the interval encapsulating zero. In practice, a skip-list is a randomized data structure where each vertex appears in a higher layer with probability $\frac{1}{2}$ to support insertion efficiently.
@@
We follow the same lazy representation as above. We store $(d_{j}^{k},s_{j}^{k})=(v_{j+1}^{k}-v_{j}^{k},s_{j}^{k})$ for edge $j$ at layer $k$. We also add a precursor vertex at the beginning of each layer. This vertex contains the smallest possible value of $\{v_{j}\}$ as serves as a boundary point. To construct $\mathsf{a}^{\prime}_{i}(\cdot)$ we update $\alpha_{0}\leftarrow\alpha_{0}+v_{0}-y_{i+1}$. Then, we follow the search procedure above until reaching the lowest layer where we calculate $z_{i}$ by solving the same linear equation. During insertion of vertex $(z_{i}-\lambda_{i},0)$, we flip a coin to decide whether it should appear in a higher layer. Note that along the searching path, we have calculated $(v_{j}^{k},\alpha_{j}^{k})$ and $(v_{j+1}^{k},\alpha_{j+1}^{k})$ per layer which contains the locus of the zero crossing. Thus, whenever a vertex should appear in a prior (higher) layer, we use the corresponding values for updating the connecting edges. We perform an equivalent update for $(z_{i}+\lambda_{i},0)$. Last, we update the $v_{0}\leftarrow v_{0}-\lambda_{i}$. The end result is an $\mathcal{O}(n\log{n})$ amortized time algorithm.

### v2

@@ 6 High-Order Variational Penalties
To illustrate the power of the convolution-based representation, we provide a specific derivation for variational penalties which correspond to second order finite differences. To do so we simply need to set $\boldsymbol{\bar{\tau}}=(2,-1)$. Since $f_{i}(\cdot)$ depends only on $x_{i-1}$ and $x_{i}$, we can represent the subgradient sets as,
$S_{i}=\{\mathsf{a}_{i}(\mathsf{x}_{i-1},\mathsf{x}_{i})\;;\;\mathsf{a}_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i})\}~{}\mbox{ and }~{}S_{i+\frac{1}{2}}=\{\mathsf{a}_{i}^{\prime}(\mathsf{x}_{i},\mathsf{x}_{i+1})\;;\;\mathsf{a}_{i+1}^{\prime}(\mathsf{x}_{i},\mathsf{x}_{i+1})\}~{}~{}.$
Assume this holds true for $S_{i+1}~{}\mbox{ i.e. }~{}S_{i+1}=\{\mathsf{a}_{i+1}(\mathsf{x}_{i},\mathsf{x}_{i+1})\;;\;\mathsf{a}_{i+2}(\mathsf{x}_{i},\mathsf{x}_{i+1})\}$. The definition of the recursive map for $i+\frac{1}{2}$ implies that,
$\mathsf{a}_{i}^{\prime}(\mathsf{x}_{i},\mathsf{x}_{i+1})=2\mathsf{a}_{i+1}(\mathsf{x}_{i},\mathsf{x}_{i+1})-\mathsf{a}_{i+2}(\mathsf{x}_{i},\mathsf{x}_{i+1})+\mathsf{x}_{i+1}-y_{i+1}~{}\mbox{ and }~{}\mathsf{a}_{i+1}^{\prime}(\cdot)=\mathsf{a}_{i+1}(\cdot)~{}~{}.$
To obtain the recursive map $S_{i+\frac{1}{2}}\to S_{i}$, we need to extract $\mathsf{x}_{i+1}$ from the following subgradient set equation,
$2\mathsf{x}_{i}-\mathsf{x}_{i-1}-\mathsf{x}_{i+1}\in\partial g_{i}^{*}(\mathsf{a}_{i}^{\prime}(\mathsf{x}_{i},\mathsf{x}_{i+1}))~{}~{}.$
@@
$T_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i})=\Big{[}2\mathsf{x}_{i}-\mathsf{x}_{i-1}\Big{]}_{z_{i+1}^{-}(\mathsf{x}_{i})}^{z_{i+1}^{+}(\mathsf{x}_{i})}~{}~{},$
where $z_{i+1}^{\pm}(\mathsf{x}_{i})$ denotes the boundary values such that $\mathsf{a}_{i}^{\prime}(\mathsf{x}_{i},z_{i+1}^{\pm}(\mathsf{x}_{i}))=\pm 1$. Using this form of $T$ back in $\mathsf{x}_{i+1}$, we obtain
@@
$\mathsf{a}_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i})=\mathsf{a}_{i+1}^{\prime}(\mathsf{x}_{i},T_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i}))~{}~{}.$
@@
Here, $\mathsf{a}_{i}(\cdot)$ is a piecewise linear manifold in $\mathbb{R}^{2}$. We store triplets of the form $\{v_{j}^{1},v_{j}^{2},\alpha_{j}\}$ for vertices of each facet defining $\mathsf{a}_{i}$. We conjecture that the number of facets grows as $\mathcal{O}(n^{2})$ but leave formal analysis of runtime as future research. We empirically observed that the description of $\mathsf{a}_{i}$ is typically linear in the length of the input sequence. Alas, a disadvantage of the exact subgradient following described here is the need to maintain a data structure for high dimensional objects. We describe an alternative approximation yielding a simple iterative procedure for the problem in Appendix D.
@@
We next demonstrate the effectiveness of exact subgradient following over the approximation-based method. To do so, we generated synthetic data of varying sequence length similar to the previous section. We sampled $(x_{i+2}-x_{i+1})-(x_{i+1}-x_{i})$ from $\mathcal{D}$ and then constructed $\mathbf{y}$ using the optimality conditions. Throughout the experiment, the sparsity level was set to $50\%$. In each experiment, we ran the approximation-based algorithm for $100$ iterations. The results are summarized in Figure 8. The advantage of exact subgradient following over approximation both in terms of runtime and optimality is apparent. The exact subgradient following weakly depends on the problem size whereas the quality of the approximate solution as well as run time quickly deteriorates even for short sequences.
7 Empirical Study
Prior to recapping the main results and concluding, we present in this section an empirical study that underscores the potential of the high-order Untangled Lariat. The experiments do not pit one version of algorithm versus another and claim superiority of a particular version, on the contrary. We describe experiments with three different financial tickers of different temporal characteristics in order to exhibit the merits of different versions in the light of different stochastic settings. The three datasets that we experimented with are SPX, VIX, and SFXRSA which we describe shortly below.
Figure 9: Comparison of Fused Lasso and Untangled Lariat on financial indices.
SPX is the ticker symbol for the Standard and Poor’s 500, a stock market index tracking the stock performance of 500 of the largest companies listed on stock exchanges in the United States. It is one of the most commonly followed equity indices and includes approximately 80% of the total market capitalization of US public companies. SPX is the case study of many research papers and has garnered numerous stock market models.
VIX is the ticker symbol for the Chicago Board Options Exchange’s CBOE Volatility Index, a popular measure of the stock market’s expectation of volatility based on S&P 500 index options. It is calculated and disseminated on a real-time basis by the CBOE, and is often referred to as the fear index or fear gauge. The VIX index tends to oscillate between periods of low volatility with a small magnitude of it and changes with spike-like transient periods. The VIX is rather difficult to model.
SFXRSA is a S&P Case-Shiller Home Price Index that measures the price level of existing single-family homes in San Francisco. The index reflects the average change in home prices in terms of percentage changes in housing market prices given a constant level of quality. Changes in the types and sizes of houses or changes in the physical characteristics of houses are specifically excluded from the calculations to avoid incorrectly affecting the index value. It is thus a slowly varying signal with some seasonal effect.
The three particular variants that we examined were the original Fused Lasso, and Untangled Lariat with a second order penalty equivalent for penalizing the total variation of the discrete second derivative. For comparison we also ran the Fused Lasso on the sequence of differences $y_{i+1}-y_{i}$. The latter mimics the $2$’nd order Untangled Lariat. The discrete derivative suppresses the low frequency components of the original sequence, obviating completely the zero frequency component. We tuned $\lambda$ for the three variants below so that their residual error of the entire series of each financial index is the same. Thus, the different behavior of each filtering approach reflects their characteristics and merits. As written above, the goal of the comparison is by no means to show superiority or a universal advantage of the nascent Untangled Lariat over the Fused Lasso or other methods mentioned in this paper. Indeed, there are settings in which the Fused Lasso would constitute the most viable data analysis tool, for instance during prolonged periods of economic stagnation.
The results for the three financial indices are depicted on the left column of Fig. 9. The three procedures seem to capture the overall tendency of the three different indices. The approximated indices generated by Fused Lasso are, as one would expect, piecewise constant. Overall, the approximation does not seem to provide an aesthetic refinement of the original data. We would henceforth more closely examine solely the performance of the Fused Lasso on the sequence of difference and the Untangled Lariat on the original data. Upon a first examination it is apparent that the Fused Lasso on the difference sequences is capable of capturing transient periods during which there is typically a sharp upturn followed by a similar or less pronounced downturn. However, there also seems to be periods of systematic bias within the approximated sequences. Since the finite derivatives of the original indices eliminates zero frequency effects, the biased estimates do not affect the overall error.
To further qualitatively examine the results, we provide on the right column of Figure 9 zoomed-in snippets of an upturn-downturn period. These snippets are provided for visualization purposes and were not handpicked. All three variants were provided the entire sequence, not the particular snippets we visualize. Traversing the figures from top hand-side to bottom hand-side, for SPX we see that the Fused Lasso’s approximation is over-fragmented and does not provide a faithful approximation of the up-then-down characteristics. The Fused Lasso on the sequence of price changes does capture nicely the temporal wedge shape but it oversubscribes to small changes on occasion. The $2$’nd order variant of Untangled Lariat nicely captures the wedge characteristics with about $4$ piecewise linear segments.
By construction and definition of the VIX index, it is more volatile than SPX. In the middle row we examine a snippet of the VIX index during a morally up-down period albeit with further nuanced changes that could be significant for a financial data analyst. The approximation of the Fused Lasso on the price changes in comparison with $2$’nd order Untangled Lariat underscores a natural trade-off. The latter approximates VIX by morally 4 linear segments, a period without a substantial change that is approximated as a sequence of fixed values, followed by a single segment of upturn, concluding with 3 segments of linear decrease in the index. It thus faithfully captures the overall trend. In contrast, the Fused Lasso on price changes still captures the nature of the sub-sequence is more fragmented and sensitive to local changes albeit resulting with the same approximation error as the Untangled Lariat due to the respective choices of $\lambda$. It thus provides a more pinpointed tool during high volatility periods that are absent from the overall market trend. Lastly, for the real-estate index SFXRSA, all three variants seem to capture quite faithfully the overall behavior of the slowly changing real-estate market in San Francisco.
We would like to note that we can possibly gain the best of all approaches with a further generalized version of Untangled Lariat with composite variational penalty consisting of two terms such as,
$\lambda_{1}|x_{i+1}-x_{i}|+\lambda_{2}|x_{i-1}-2x_{i}+x_{i+1}|~{}~{}.$
We leave the derivation and analysis of Untangled Lariat with composite variational penalties to future research.
8 Concluding Remarks
This work (subjectively) makes the following novel contributions:
  • Provides a unified approach for sequence approximation w.r.t. Bregman divergences with general total variation penalties.
  • Describes and analyzes a novel subgradient following procedure, Untangled Lariat, with succinct code that can be readily used with existing and new variation penalties.
  • Untangled Lariat entertains the same time complexity as known algorithms, each of which designed (by different authors) for a concrete variational penalty.
  • Introduces and analyzes a multivariate generalization where each element of the sequence is a vector. The base case of squared $2$-norm as the variational penalty of two consecutive vectors bears a closed form for subgradient following.
  • Derives iterative algorithms for sparse multivariate total variation using the $2$-norm or the $\infty$-norm by employing the norm-squared version as surrogate.
  • Describes and analyzes nascent high-order Untangled Lariat with poly-time subgradient following.
@@ 8 Concluding Remarks
$\lambda_{i}^{\mathsf{R}}\big{[}x_{i+1}-x_{i}-\epsilon_{i}^{\mathsf{R}}\big{]}_{+}\,+\,\lambda_{i}^{\mathsf{L}}\big{[}x_{i}-x_{i+1}-\epsilon_{i}^{\mathsf{L}}\big{]}_{+}~{}~{}.$
By choosing $\lambda_{i}^{\mathsf{R}}=\lambda_{i}^{\mathsf{L}}=\lambda_{i}$ and $\epsilon_{i}^{\mathsf{R}}=\epsilon_{i}^{\mathsf{L}}=0$ we obtain the fused lasso’s penalty. To encompass this version we simply need to compute $\partial g_{i}^{*}$ from which we derive a new mapping $\mathsf{x}\mapsto\mathsf{a}_{i}(\mathsf{x})$. The subgradient procedure uses a similar form for the mapping given in Eq. (7) though the introduction of insensitivity regions to the left and/or the right of the loci where $x_{i}=x_{i+1}$ would introduce two additional inflection points. The resulting procedure for constructing the subgradient would require at most $\mathcal{O}(n\log n)$ time.
We limited our discussion to sequential penalties which can described as a chain of local connectivities $x_{1}\leftrightarrow x_{2}\leftrightarrow\ldots x_{i}\leftrightarrow x_{i+1}\ldots$ forming a left-to-right dependency graph. Kolmogorov et. al [18] examined a setting where the total variation penalties form a tree. Denote the variation dependency graph by $G=(V,E)$ where $V=[n]$ and $(i,j)\in E$ if there exists a penalty term for the variables $x_{i}$ and $x_{j}$. Using the graph representation, the penalized problem takes the form of,
@@
$U$ $=$ $\max_{\|\boldsymbol{\alpha}\|_{\infty}\leq 1}\,\min_{{\mathbf{x}}}\>\frac{1}{2}\sum_{i=1}^{n}(x_{i}-y_{i})^{2}+\sum_{(i,j)\in E}\alpha_{i,j}(x_{i}-x_{j})~{}~{}.$
@@ Acknowledgments
We would like to thank Tomer Koren and Yishay Mansour for fruitful discussions. The short discussion above is motivated by conversations with the late Prof. Tali Tishby who is missed dearly. This manuscript was written without the usage of AIML tools.
Appendix A Improved Barrier Lariat using Skiplists
@@
We store each $\mathsf{a}_{i}(\cdot)$ as a linked-list data structure of implicit change points (vertices) and explicit linear intervals connecting two consecutive vertices. A vertex $(v_{j},\mathsf{a}_{i}(v_{j}))=(v_{j},\alpha_{j})$ represents a point where $\mathsf{a}_{i}(\cdot)$ changes slope. An edge $(d_{j},s_{j})$ stores the length and slope of the line segment between $v_{j}$ and $v_{j+1}$, thus we have,
$d_{j}=v_{j+1}-v_{j}\,;\,s_{j}=(\alpha_{j+1}-\alpha_{j})/d_{j}~{}~{}.$
@@
This representation allows lazy evaluations in a similar fashion to the fused lasso. The Kvetsh step does not change the lengths of intervals except for the new segment encapsulating the zero crossing. Since the length of all other line segments remain intact this edge-based representation facilitates an efficient implementation.
By construction of $\mathsf{a}_{i+1}$ we know that $\exists j$ for which $v_{j}=z_{i+1}-\lambda_{i+1}$ and $\mathsf{a}_{i+1}(z_{i+1}-\lambda_{i+1})=0$. This implies that in an explicit representation of $\mathsf{a}^{\prime}_{i}(\cdot)$ we would need to update vertex $j$,
$(z_{i+1}-\lambda_{i+1},0)~{}~{}\mapsto~{}~{}(z_{i+1}-\lambda_{i+1},z_{i+1}-\lambda_{i+1}-y_{i+1})~{}~{}.$
@@
In the explicit view of the lazy representation of the ensuing edge, the slope $s_{j}$ of the segment $[v_{j},v_{j+1}]$ becomes $s_{j}+n-i$. We next find the zero crossing point of $\mathsf{a}^{\prime}_{i}(\cdot)$ by traversing the list of edges towards zero and locating the index $k$ for which $\alpha_{k}\leq 0$ and $\alpha_{k+1}>0$. In case such an index does not exist the zero of $\mathsf{a}^{\prime}_{i}(\cdot)$ is right of its rightmost vertex. Once we identify the enclosing interval, we calculate $z_{i}$ by solving a linear equation from $(v_{k},\alpha_{k})$ and $(v_{k+1},\alpha_{k+1})$. We next Kvetsh the list by inserting two vertices, $(z_{i}-\lambda_{i},0)$ and $(z_{i}+\lambda_{i},0)$, between $k$ and $k+1$, and update $d_{k}$ accordingly. This sequential traversal of the list in the worst-case would require linear time and thus total run time would remain intact at $\mathcal{O}(n^{2})$.
@@
In order to employ a more efficient procedure for finding zero crossings, we make use of a skip-list data structure [20]. A skip-list is an extension of a linked list with multiple layers and with which the expected search time is $\mathcal{O}(\log{n})$. The skip-list has ideally in our setting $\mathcal{O}(\log{n})$ layers where layer $k$ contains all vertices whose indices modulo $2^{k}$ are zero. The search procedure starts from the top layer and proceeds downwards once we reach the interval encapsulating zero. In practice, a skip-list is a randomized data structure where each vertex appears in a higher layer with probability of $\frac{1}{2}$ to support efficient insertions.
@@
We follow the same lazy representation as above. We store $(d_{j}^{k},s_{j}^{k})=(v_{j+1}^{k}-v_{j}^{k},s_{j}^{k})$ for edge $j$ at layer $k$. We also add a precursor vertex at the beginning of each layer. This vertex contains the smallest possible value of $\{v_{j}\}$ and serves as a boundary point. To construct $\mathsf{a}^{\prime}_{i}(\cdot)$ we update $\alpha_{0}\leftarrow\alpha_{0}+v_{0}-y_{i+1}$. Then, we follow the search procedure above until reaching the bottom layer. We calculate $z_{i}$ at the bottom layer by solving the same linear equation as above. During the insertion of vertex $(z_{i}-\lambda_{i},0)$, we flip a coin to decide whether it should appear in a higher layer. Note that along the searching path, we have calculated $(v_{j}^{k},\alpha_{j}^{k})$ and $(v_{j+1}^{k},\alpha_{j+1}^{k})$ per layer which contains the locus of the zero crossing. Thus, whenever a vertex should appear in a prior (higher) layer, we use the corresponding values for updating the connecting edges. We perform an equivalent update for $(z_{i}+\lambda_{i},0)$. Last, we update the $v_{0}\leftarrow v_{0}-\lambda_{i}$. The end result is an $\mathcal{O}(n\log{n})$ amortized time algorithm.